OpenStudy (anonymous):

Solve and check your answer. 2/ 1-x = 5/ 1+x

OpenStudy (tkhunny):

First, write it correctly. 2/(1-x) = 5/(1+x) maybe?

OpenStudy (anonymous):

That is how the instructor wrote it. =l

OpenStudy (anonymous):

2/ 1-x = 5/ 1+x cross multiplication we get; 2(1+x)=5(1-x) 2x+2=5-5x 2x+5x=5-2 7x=3 x=3/7 on checking: put x=3/7 on above equation u get; 2/1-(3/7)=5/1+(3/7) 7/2=7/2 which is true.. therefore x=3/7 is the reqd answer..

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome

OpenStudy (tkhunny):

@PECKCE "That is how the instructor wrote it. =l " I sincerely hope not. Was the instructir writing in an on-line environment? If this is the case, then this is your big chance to correct the instructor who seems to have forgotten the Order of Operations. Those parentheses are NOT optional.